# A search for super-imposed oscillations to the primordial power spectrum in Planck and SPT-3G 2018 data

### v1

@@ IV Conclusions
We have searched for super-imposed oscillations linearly or logarithmically spaced in Fourier wavenumbers in Planck and SPT-3G 2018 temperature and polarization data. SPT-3G angular resolution and sensitivity in polarization can test primordial features at higher multipoles beyond Planck measurements. SPT-3G 2018 temperature and polarization data [35] lead to $\Lambda$CDM parameters very consistent with Planck and can be therefore used in combination with Planck to provide a precision test for models beyond $\Lambda$CDM such as those including features in the primordial power spectrum. See [35, 44, 45] for other studies of models beyond $\Lambda$CDM with SPT-3G data.
@@
In this analysis we use SPT-3G public (binned) data. We therefore use also Planck binned 2018 data, differently from the search for features performed in Ref. [1]. We use polychord [39] as a sampler given the multimodality of the resulting posterior distributions and we allow for variation of the foreground and nuisance parameters.
@@
For Planck, we find results consistent with those obtained in [1] for oscillations with a constant amplitude, which produce a quite similar improvement in the fit compared to $\Lambda$CDM. When these oscillations have a Gaussian modulation, linear and logarithmic oscillations produce $\sim-13$ and $\sim-11$ as improvement in $\Delta\chi^{2}$, respectively.
@@
We have tested these models with primordial oscillations with SPT-3G 2018 data for the first time. For SPT-3G 2018 data, we have found that all the models studied here improve the fit compared to $\Lambda$CDM, although are not statistically preferred over the $\Lambda$CDM. As happens for $\Lambda$CDM cosmology [35], SPT-3G is less constraining than Planck 2018 data. Our analysis shows that the frequencies preferred by SPT-3G 2018 data do not coincide with those preferred by Planck 2018 data when the amplitude of the super-imposed oscillations is constant. However, for a Gaussian modulated amplitude, we find prominent peaks at $\log_{10}\omega_{\rm log}\sim 1.6$ and at $\log_{10}\omega_{\rm lin}\sim 1.1$. The peak in the frequency for modulated linear oscillations found in SPT-3G 2018 lies very close to the found in Planck 2018 data. In particular, we find that super-imposed logarithmic oscillations lead to $\Delta\chi^{2}\sim-12$, and when Gaussian modulated to $\Delta\chi^{2}\sim-14$. These oscillations do not correspond exactly to oscillations in SPT-3G EE data found by Gaussian regression in [42]  [^4] Oscillations in $C_{\ell}^{EE}$ are not necessarily fit by super-imposed oscillations to the curvature power spectrum which are also imprinted in $C_{\ell}^{TT}$..
@@
We have therefore constrained these models with the combination of Planck and SPT-3G data for the first time. Since a covariance between Planck and SPT-3G data is not available, we consider the main combined data set in which we cut Planck (SPT) above (below) the multipoles at which the Planck S/N is reached by SPT as stated in [35] and described in section II. This conservative procedure which enforces just one measurement for each multipole bin should overcome the issue of potentially double counting information with a full Planck and SPT-3G combination without a covariance matrix between the two experiments. By adopting this combination, we find that SPT-3G $-\ell$ polarization data play an important role for all the models studied.
@@
We find that superimposed oscillations with a constant amplitude do not improve the fit and are tighter constrained by Planck+SPT-3G 2018 compared to Planck 2018 data. When the ranges of parameters which provide a better fit independently to Planck and SPT-3G data overlap, we find a larger $\Delta\chi^{2}$ in the combined Planck SPT-3G data set, up to the maximum $\Delta\chi^{2}\sim-17$ for modulated logarithmic oscillations.
@@
Our findings could be extended to higher frequencies with data which are either unbinned or have a smaller binning and will be further tested with upcoming CMB temperature and polarization measurements at high multipoles by ongoing and future ground experiments, such as ACT, SPT, Simons Observatory [47] and CMB-S4 [48].
Acknowledgements.
@@
The authors acknowledge the use of computational resources at the Institute of Mathematical Science’s High Performance Computing facility [Nandadevi] and of CNAF HPC cluster in Bologna. FF and DKH would like to thank Karim Benabed and Silvia Galli for discussions on SPT-3G 2018 likelihood. AA is supported by an appointment to the JRG Program at the Asia Pacific Center for Theoretical Physics through the Science and Technology Promotion Fund and Lottery Fund of the Korean Government, and was also supported by the Korean Local Governments in Gyeongsangbuk-do Province and Pohang City.AA also acknowledges support from the NRF of Korea (Grant No. NRF-2022R1F1A1061590) funded by the Korean Government (MSIT). AA, FF, DKH acknowledge travel support through the India-Italy ”RELIC - Reconstructing Early and Late events In Cosmology” mobility program. FF would like to thank IMSc for warm hospitality for the December 2022 visit when this work started. FF, DP acknowledge financial support from the contract by the agreement n. 2020-9-HH.0 ASI-UniRM2 “Partecipazione italiana alla fase A della missione LiteBIRD”. FF, DKH, DP acknowledge financial support from Progetti di Astrofisica Fondamentale INAF 2023. DKH would like to acknowledge the support from the Indo-French Centre for the Promotion of Advanced Research – CEFIPRA grant no. 6704-4. AS would like to acknowledge the support by National Research Foundation of Korea NRF-2021M3F7A1082056, and the support of the Korea Institute for Advanced Study (KIAS) grant funded by the government of Korea.

### v2

@@ 4 Conclusions
We have searched for super-imposed oscillations linearly or logarithmically spaced in Fourier wavenumbers in Planck and SPT-3G 2018 temperature and polarization data. We study these super-imposed oscillations either with a constant amplitudes or amplitudes modulated with a Gaussian envelope; to our knowledge it is the first time that the modulated logarithmic oscillation is tested with current CMB data. SPT-3G, with better angular resolution and sensitivity can test primordial features at multipoles beyond the $Planck$ resolution. This is particularly advantageous for features that are at the edge of the $Planck$ angular resolution as the one mimicking the $A_{\rm L}$ effect. Cosmological parameters estimated from SPT-3G 2018 temperature and polarization data [48] in power law $\Lambda$CDM model are completelely consistent with $Planck$, and therefore, can be used in combination with $Planck$ to provide a precision test for models beyond the power law model such as those including features in the primordial power spectrum. See [48, 56, 57] for other studies of models beyond power law with SPT-3G data.
@@
In this analysis, we use SPT-3G public (binned) data. We therefore use also $Planck$ binned 2018 data, differently from the search for features performed in Ref. [1], which used unbinned data but fixed nuisance and foreground parameters to their baseline values. We use PolyChord [50] as a sampler given the multimodal nature of the resulting posterior distributions and we allow for variation of the foreground and nuisance parameters.
@@
For $Planck$, we find results consistent with those obtained in [1] for oscillations with a constant amplitude. When these oscillations have a Gaussian modulation, linear and logarithmic oscillations produce $\sim-11.8$ and $\sim-10.0$ as improvement in $Planck$ 2018 $\Delta\chi^{2}$, respectively.
@@
We have tested these models with primordial oscillations with SPT-3G 2018 data for the first time. For SPT-3G 2018 data, we have found that all the models studied here improve the fit compared to the power law model, although they are not statistically preferred over the power-law power spectrum. As expected, for $\Lambda$CDM+power law cosmology [48], SPT-3G is less constraining than $Planck$ 2018 data. Our analysis shows that the frequencies preferred by SPT-3G 2018 data do not coincide with those preferred by $Planck$ 2018 data when the amplitude of the super-imposed oscillations is constant. However, for a Gaussian modulated amplitude, we find a prominent peak at $\log_{10}\omega_{\rm log}\sim 1.6$ and a broader peak at $\log_{10}\omega_{\rm lin}\gtrsim 1.1$. In particular, we find that super-imposed logarithmic oscillations lead to $\Delta\chi^{2}\sim-12$. These oscillations do not correspond exactly, but are related to certain features in SPT-3G EE data found by Gaussian regression in [55]  [^4] Oscillations in $C_{\ell}^{EE}$ are not necessarily fit by super-imposed oscillations to the curvature power spectrum which are also imprinted in $C_{\ell}^{TT}$..
@@
We have therefore constrained these models with the combination of $Planck$ and SPT-3G data for the first time. Since a covariance between $Planck$ and SPT-3G data is not available, we consider the main combined data set in which we cut $Planck$ (SPT) above (below) the multipoles at which the $Planck$ S/N is reached by SPT as stated in [48] and described in section 2. This conservative procedure which enforces just one measurement for each multipole bin should overcome the issue of potentially double counting information with a full $Planck$ and SPT-3G combination without a covariance matrix between the two experiments. By adopting this combination, we find that SPT-3G $-\ell$ polarization data play an important role for all the models studied.
@@
We find that superimposed oscillations with a constant amplitude do not improve the fit and are tighter constrained by $Planck$+SPT-3G 2018 compared to $Planck$ 2018 data. When the ranges of parameters which provide a better fit independently to $Planck$ and SPT-3G data overlap, we find a larger $\Delta\chi^{2}$ in the combined $Planck$ SPT-3G data set, $\Delta\chi^{2}\sim-17.5$ (-14.7) for modulated logarithmic (linear) oscillations. It is interesting to note that SPT-3G data do not prefer $A_{L}$ larger than 1 [48], but do not disfavour the modulated linear oscillation template which mimics the $A_{L}$ effect [1]. This different response of SPT-3G data to different extensions of the power law model might be due to the number of extra parameters, 1 for $A_{L}$ and 5 for the modulated oscillation template.
@@
Our findings could be extended to higher frequencies with data which have a finer binning and will be further tested by upcoming CMB temperature and polarization measurements at high multipoles by ongoing and future ground experiments, such as ACT, SPT, Simons Observatory [58] and CMB-S4 [59].
Acknowledgments
@@
The authors acknowledge the use of computational resources at the Institute of Mathematical Science’s High Performance Computing facility [Nandadevi] and of CNAF HPC cluster in Bologna. We would like to thank Jan Hamann for useful comments on the manuscript. FF and DKH would like to thank Karim Benabed and Silvia Galli for discussions on SPT-3G 2018 likelihood. AA is supported by an appointment to the JRG Program at the Asia Pacific Center for Theoretical Physics through the Science and Technology Promotion Fund and Lottery Fund of the Korean Government, and was also supported by the Korean Local Governments in Gyeongsangbuk-do Province and Pohang City. AA also acknowledges support from the NRF of Korea (Grant No. NRF-2022R1F1A1061590) funded by the Korean Government (MSIT). AA, FF, DKH acknowledge travel support through the India-Italy "RELIC - Reconstructing Early and Late events In Cosmology" mobility program. FF would like to thank IMSc for warm hospitality for the December 2022 visit when this work started. FF, DP acknowledge financial support from the contract by the agreement n. 2020-9-HH.0 ASI-UniRM2 “Partecipazione italiana alla fase A della missione LiteBIRD". FF, DKH, DP acknowledge financial support from Progetti di Astrofisica Fondamentale INAF 2023. DKH would like to acknowledge the support from the Indo-French Centre for the Promotion of Advanced Research – CEFIPRA grant no. 6704-4. AS would like to acknowledge the support by National Research Foundation of Korea NRF-2021M3F7A1082056, and the support of the Korea Institute for Advanced Study (KIAS) grant funded by the government of Korea.